OpenStudy (anonymous):

Find the value of a so that the period(T) of the function y=cos^2[(a^2+2a-28)*x)] would be T=PI/20

OpenStudy (experimentx):

which is the variable x or a??

OpenStudy (anonymous):

I'm not actually sure, but I think it might be x.

OpenStudy (anonymous):

a is supposed to be a parameter.

OpenStudy (experimentx):

i don't think you will get time period so easily ... if there is a general term a if x is a variable.

OpenStudy (anonymous):

In which case the problem would make sense the most?

OpenStudy (experimentx):

sorry, i didn't read the question properly

OpenStudy (anonymous):

Oh, alright.

OpenStudy (experimentx):

(a^2+2a-28) = pi/20 or, (a+1)^2 = pi/20 + 29 now solve for a

OpenStudy (anonymous):

And that's it?

OpenStudy (experimentx):

yes, http://www.wolframalpha.com/input/?i=y%3Dcos%5E2%5B10*x%29%5D+

OpenStudy (anonymous):

Alright, thanks!

OpenStudy (asnaseer):

I /think/ is you have a trig function of the form y = cos(Ax), then it's period is given by \(\frac{2\pi}{A}\). so first you need to convert the \(\cos^2\) function into one which does not involve powers. I believe you can use this formulae:\[\cos^2(x)=\frac{\cos(2x)+1}{2}\]

OpenStudy (asnaseer):

replace "is" with "if" after "/think/" above

OpenStudy (asnaseer):

so you would end up with:\[\cos^2((a^2+2a-28)x)=\frac{\cos(2(a^2+2a-28)x)+1}{2}\]

OpenStudy (asnaseer):

and to get a period of \(\frac{\pi}{20}\) you need to solve:\[\frac{2\pi}{2(a^2+2a-28)}=\frac{2\pi}{20}\]

OpenStudy (anonymous):

And the a value I get is the answer?

OpenStudy (asnaseer):

yes - you should get two values since this is a quadratic equation

OpenStudy (asnaseer):

sorry - I made a mistake up there, the right-hand-side should be \(\frac{\pi}{20}\)

OpenStudy (asnaseer):

not \(\frac{2\pi}{20}\)

OpenStudy (anonymous):

Oh, alright. Thanks!

OpenStudy (asnaseer):

yw

OpenStudy (experimentx):

period was already changed by half when you made it 2x

OpenStudy (asnaseer):

no - I just rewrote cos^2(x) in terms of cos(2x)

OpenStudy (asnaseer):

the resulting equation is still the same equation, so we still need it to have the same period

OpenStudy (experimentx):

quite not ... the period of [cos(x)]^2 will be 2pi while period of cos(2x) will be pi it seems that way. http://www.wolframalpha.com/input/?i=y%3Dcos%5E2%5B10*x%29%5D+ the time period of this expression still seem to be 2pi/x http://www.wolframalpha.com/input/?i=y%3D+%28cos%5B20*x%29%5D+%2B1%29%2F2++ while isolating this expression will change the time period by pi/x http://www.wolframalpha.com/input/?i=y%3Dcos+%5B20*x%29%5D+

OpenStudy (asnaseer):

experimentX: do you agree that:\[cos^2(x)=\frac{cos(2x)+1}{2}\]

OpenStudy (experimentx):

yeah i agree, the both expression on the right hand side and left hand side will have the same time period,

OpenStudy (asnaseer):

so what are you trying to say then? I don't understand?